OpenStudy (anonymous):

a +2 RIDDLE OF THE DAY A car travels along a straight line for the first half time with a speed of 40 km/h and the second half of 60 km/h find the average speed of the car.

OpenStudy (anonymous):

@Mimi_x3 @mathslover @ParthKohli @pr3dat0r

OpenStudy (anonymous):

@TheViper wanna try

OpenStudy (lgbasallote):

let d = distance \[\frac{60t + 40t}{2t} \implies \frac{100t}{2t} = 50m/s\]

OpenStudy (lgbasallote):

let t= time not d = distance

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

lol @Rohangrr do you even know the answers to these

OpenStudy (anonymous):

yes just posting having a lot of time

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol^2

OpenStudy (lgbasallote):

\[Speed_{ave} = \frac{\text{total distance }}{\text{total time}}\]

OpenStudy (lgbasallote):

total distance = 60t + 40t total time = t + t =2t

OpenStudy (anonymous):

Technically, speed is\[s=|v_{avg}|=\left|\frac{\Delta x}{\Delta t}\right|\]irrelevant for this problem, though.